# ARMAr-LASSO: Mitigating the Impact of Predictor Serial Correlation on the LASSO

### v2

@@ 3 The ARMAr-LASSO
Model (4) is the linear projection of $y_{t}$ on $\mathbf{w}_{t}=(u_{1,t},\dots,u_{n,t},y_{t-1},\dots,y_{t-p_{y}})^{\prime}$, which contains $n$ ARMA residuals and $p_{y}$ lagged values of the response. $\boldsymbol{\beta}^{*}=(\boldsymbol{\alpha}^{*^{\prime}},\phi_{y,1},\dots,\phi_{y,p_{y}})^{\prime}$ represents the corresponding best linear projection coefficients and $v_{t}$ is the error term, which is unlikely to be a white noise. It should be noted that the choice of $p_{y}$ is arbitrary and that some lags will be relevant while others will not. The relevant lags will be directly selected using LASSO. Model (4) is misspecified, in the sense that it does not correspond to the true data generating process (DGP) for the response, but it is similar in spirit to the factor filter used in the literature to mitigate multicollinearity (Fan et al., 2020). The idea behind model (4) is to leverage the serial independence of the $u$ terms, thereby avoiding the risk of sizeable spurious correlation. However, the $u$ terms alone may explain only a small portion of the variance of $y_{t}$, particularly in situations with high persistence. This is why we introduce the response lags as additional predictors; these amplify the signal in our model and consequently improve the forecast of $y_{t}$. We list some important facts that capture how misspecification affects coefficient estimation and feature selection.
@@
(common AR(1) restriction). Suppose both predictors and error terms in model (1) admit an AR(1) representation; that is, $x_{i,t}=\phi x_{i,t-1}+u_{i,t}$ and $\varepsilon_{t}=\phi\varepsilon_{t-1}+\omega_{t}$. In this case $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t}+\varepsilon_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}(\phi x_{i,t-1}+u_{i,t})+\phi\varepsilon_{t-1}+\omega_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}u_{i,t}+\phi y_{t-1}+\omega_{t}.$ Thus, under the common AR(1) restriction (also known as common factor restriction,  Mizon, 1995), the working model (4) is equivalent to the true model (1) because of the decomposition of the AR(1) processes $\{\mathbf{x}_{t}\}$ and $\{\varepsilon_{t}\}$.
@@
The working model (4) coincides with the true model (1) under a common AR($p$) restriction; that is, when $x_{i,t}=\sum_{l=1}^{p}\phi_{l}x_{i,t-l}+u_{i,t}$ and $\varepsilon_{t}=\sum_{l=1}^{p}\phi_{l}\varepsilon_{t-l}+\omega_{t}$. In fact, it is easy to show that $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t}+\varepsilon_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}u_{i,t}+\sum_{l=1}^{p}\phi_{l}y_{t-l}+\omega_{t}$ for any autoregressive order $p$. Moreover, in this case $v_{t}=\omega_{t}$ and $E(v_{t}|\mathbf{w}_{t})=0$ – so we have unbiasedness and consistency also for the coefficients of the lags of $y_{t}$.
@@
(common factor). Suppose $x_{i,t}=\lambda_{i}f_{t}+z_{t},\ f_{t}=\phi_{f}f_{t-1}+\delta_{t},\ z_{i,t}=\phi_{i}z_{i,t-1}+\eta_{i,t}$ and $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t},$ where $\delta_{t},\eta_{it},\omega_{t}\sim i.i.d\ N(0,1).$ In this case, any $x_{it}$ is a sum of two independent AR(1) processes and, therefore, $x_{it}\sim ARMA(2,1)$ (Granger and Morris, 1976). Again, by Assumption 1, we have $E(v_{t}|\mathbf{u}_{t})=0$ and $E(v_{t}|y_{t-1})\neq 0$.
@@
Let $\mathbf{q}_{t}=(\mathbf{w}_{t}^{\prime},v_{t})^{\prime}$. To derive theoretical results for ARMAr-LASSO, we rely on the fact that, due to Assumption 1, $\mathbf{q}_{t}$ depends almost entirely on the “near epoch” of its shock. In particular, it is characterized as near-epoch dependent (NED) (refer to  Davidson, 1994 ch. 17 and  Adamek et al., 2023 for details). NED is a very popular tool for modelling dependence in econometrics. It allows for cases where a variable’s behaviour is primarily governed by the recent history of explanatory variables or shock processes, potentially assumed to be mixing. Davidson (1994) shows that even if a variable is not mixing, its reliance on the near epoch of its shocks makes it suitable for applying limit theorems, particularly the mixingale property (see Supplement B.4 for details). The NED framework accommodates a wide range of models, including those that are misspecified as our working model (4). For instance, in Examples 2 and 3, $(\mathbf{w}_{t}^{\prime},v_{t})$ have a moving average representation with geometrically decaying coefficients, and are thus NED on $(\mathbf{u}_{t}^{\prime},\omega_{t})$ and $(\delta_{t},\boldsymbol{\eta}_{t}^{\prime},\omega_{t})$, respectively.
3.1 Asymptotic Results
@@
This section is devoted to the asymptotic behaviour and feature selection consistency of the LASSO applied to working model (4), within the classical setting with $n$ fixed and $T\rightarrow\infty$. We will extend some known results to our context to demonstrate that the working model (4) retains the usual inferential and selection consistency properties, despite being a misspecification of the true model (1). Our results build upon Theorem 2 of Fu and Knight (2000) and Theorem 1 of  Zhao and Yu (2006). We note that, when establishing asymptotic results, we can consider the $u$’s and neglect the problem of their estimation. In fact, properties of maximum likelihood estimators for the parameters of ARMA models guarantee that, on large samples, estimated ARMA residuals behave like the non-observable true ones (see Reinsel 1993, p. 117). Let $\boldsymbol{\mu}_{vy}=\mathopen{}\mathclose{{}\left(E(v_{t}y_{t-1}),\dots,E(v_{t}y_{t-p_{y}})}\right)^{\prime}$ be the mean vector and $\boldsymbol{\Gamma}_{vy}$ the $p_{y}\times p_{y}$ covariance matrix of $\mathopen{}\mathclose{{}\left(v_{t}y_{t-1},\dots,v_{t}y_{t-p_{y}}}\right)$. The following theorem provides the asymptotic behaviour of the LASSO solution.
@@
(strong irrepresentable condition (Zhao and Yu, 2006)) Assuming $\widehat{\mathbf{C}}_{w}(11)$ is invertible, $|\widehat{\mathbf{C}}_{w}(21)(\widehat{\mathbf{C}}_{w}(11))^{-1}Sign(\boldsymbol{\beta}^{*}(1))|<\boldsymbol{1}-\varphi,$ where $\varphi\in(0,1)$ and the inequality holds element-wise.
Zhao and Yu (2006) showed that Assumption 2 is sufficient and almost necessary for both estimation and sign consistencies of the LASSO. The former requires $||\widehat{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}||\stackrel{{\scriptstyle p}}{{\rightarrow}}0$, for some norm |||||$\cdot$||| | ⋅ | | (see  Fan et al., 2020). The latter requires $\underset{T\rightarrow\infty}{\text{lim}}P(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*}))=1$ and implies selection consistency; namely, $\underset{T\rightarrow\infty}{\text{lim}}P(Supp(\widehat{\boldsymbol{\beta}})=Supp(\boldsymbol{\beta}^{*}))=1$.  Zhao and Yu (2006) also provided some conditions that guarantee the strong irrepresentable condition. The following are examples of such conditions: when $|\widehat{c}_{ij}|<\frac{1}{2||\boldsymbol{\beta}^{*}||_{0}-1}$ for any $i\neq j$ (Zhao and Yu, 2006, Corollary 2); when $\widehat{c}_{ij}=\rho^{|i-j|}$ for $|\rho|<1$ (Zhao and Yu, 2006, Corollary 3); or when these conditions are block-wise satisfied (Zhao and Yu, 2006, Corollary 5). As a consequence of Fact 1 ($b$), $\widehat{\mathbf{C}}_{w}$ exhibits a block-wise structure, whereby one block encompasses the correlations between $u$’s and another block encompasses the correlations between lags of $y_{t}$. Thus, Assumption 2 is satisfied if, for instance, the bound $\frac{1}{2||\boldsymbol{\beta}^{*}||_{0}-1}$ holds for the first block and the power decay bound $\rho^{|i-j|}$ holds for the second (see also Nardi and Rinaldo, 2011). The following theorem states the selection consistency of our LASSO solution under Assumption 2.
Theorem 3:
Let Assumptions 1 and 2 hold. If $\lambda\sqrt{T}\rightarrow\lambda_{0}\geq 0$, then the solution $\widehat{\boldsymbol{\beta}}$ of (5) is such that $P\mathopen{}\mathclose{{}\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}\right)\rightarrow 1.$
Proof: see Supplement B.3
The theoretical results provided in this section show that under Assumptions 1 and 2, and as a consequence of Fact 1, ARMAr-LASSO guarantees consistent estimation, asymptotic normality, as well as consistent feature selection for the vector $\boldsymbol{\alpha}^{*}$.
3.2 Oracle Inequalities
In this section, we derive the oracle inequalities that provide bounds for the estimation and forecast errors of the ARMAr-LASSO. Here, we allow $n$ to grow as $T$ grows, that is, henceforth results are in a framework where $n=n_{T}$. We need the following two Assumptions, which bound the unconditional moments of the predictors in the true model (1), and of the predictors and errors in the working model (4).
Assumption 3:
There exist constants $b_{2}>b_{1}>2$ such that $\underset{i\leq n_{T},t\leq T}{\text{max}}E(\mathopen{}\mathclose{{}\left|x_{i,t}}\right|^{2b_{2}})\leq C$.
Assumption 4:
Consider $\mathbf{q}_{t}=(\mathbf{w}_{t}^{\prime},v_{t})^{\prime}$. There exist constants $c_{2}>c_{1}>2$ such that $\underset{i\leq n_{T}+p_{y}+1,\ t\leq T}{\text{max}}E(\mathopen{}\mathclose{{}\left|q_{i,t}}\right|^{2c_{2}})\leq C$.
@@ 3.2 Oracle Inequalities
Let Assumptions 1, 3 and 4 hold. Furthermore, assume that $T$ and $n_{T}$ are sufficiently large as to have $\lambda\geq C\mathopen{}\mathclose{{}\left(\sqrt{\text{log(T)}}}\right)^{1/c_{1}}\cfrac{(n_{T}+p_{y})^{1/c_{1}}}{\sqrt{T}}$. Then, the following inequalities hold simultaneously with probability at least $1-C\mathopen{}\mathclose{{}\left(\sqrt{\text{log(T)}}}\right)^{-1}$: ($a$) $\underset{i\leq n_{T}+p_{y},l\leq T}{\text{max}}\mathopen{}\mathclose{{}\left|\sum_{t=1}^{l}w_{i,t}v_{t}}\right|\leq\frac{T\lambda}{4}$; ($b$) $\underset{i\leq n_{T}}{\text{max}}\mathopen{}\mathclose{{}\left|\mathopen{}\mathclose{{}\left|\widehat{\boldsymbol{\gamma}_{i}}-\boldsymbol{\gamma}_{i}}\right|}\right|_{\infty}\leq C$.
Proof: see Supplement B.4.
Theorem 4 establishes that the inequalities we need for the error bound of the proposed ARMAr-LASSO estimator hold with high probability. The bounds used in the proof of Theorem 4 put implicit limits on the divergence rate of $n_{T}$ relative to $T$ (see Lemmas B.5 and B.6 in Supplement B.4). The term $\sqrt{\text{log($T$)}}$ is chosen arbitrarily as a sequence that grows slowly as $T\rightarrow\infty$. However, we can use any sequence that tends to infinity sufficiently slowly. For example, Adamek et al. (2023) use $\text{log}(\text{log}(T))$ to derive properties of the LASSO in a high-dimensional time series model under weak sparsity. Note that the importance of inequality ($b$) stems from the role of $\widehat{\boldsymbol{\gamma}_{i}}$ in the estimation of ARMA coefficients (see  Brockwell and Davis, 2002, ch. 5). We introduce an assumption on the “restricted” positive definiteness of the covariance matrix of the predictors, which allows us to generalize subsequent results to the high-dimensional framework.

### v3

@@ 3 The ARMAr-LASSO
Model (1) is the linear projection of $y_{t}$ on $\mathbf{w}_{t}=(u_{1,t},\dots,u_{n,t},y_{t-1},\dots,y_{t-p_{y}})^{\prime}$, which contains $n$ ARMA residuals and $p_{y}$ lagged values of the response. $\boldsymbol{\beta}^{*}=(\boldsymbol{\alpha}^{*^{\prime}},\phi_{y,1},\dots,\phi_{y,p_{y}})^{\prime}$ represents the corresponding best linear projection coefficients and $v_{t}$ is the error term, which is unlikely to be $i.i.d.$. It should be noted that the choice of $p_{y}$ is arbitrary and that some lags will be relevant while others will not. The relevant lags will be directly selected using LASSO. Model (1) is misspecified, in the sense that it does not correspond to the true data generating process (DGP) for the response, but it is similar in spirit to the factor filter used in the literature to mitigate multicollinearity (Fan2020). The idea behind model (1) is to leverage the serial independence of the $u$ terms, thereby avoiding the risk of sizeable spurious correlation. However, the $u$ terms alone may explain only a small portion of the variance of $y_{t}$, particularly in situations with high persistence. This is why we introduce the response lags as additional predictors; these amplify the signal in our model and consequently improve the forecast of $y_{t}$. Furthermore, the inclusion of lagged terms also helps mitigating serial correlation in the residuals.
Remark 1:
When $p_{y}=0$, ARMAr-LASSO uses only the $u_{t}$ as predictors, leaving $v_{t}$ to capture all serial correlation. Including lags of $y_{t}$ ($p_{y}>0$) allows one to model and exploit serial correlation. Under the common AR($p$) restriction, $s_{y}=p$ ensures that $v_{t}$ is a white noise; using fewer lags ($s_{y}<p$) reduces forecasting performance. Without this restriction, ARMAr-LASSO automatically selects $s_{y}$ based on the maximum AR/ARMA order of the predictors and errors. Although a formal theoretical proof of this result is beyond the scope of the present study, our numerical evidence shows that when $p_{y}$ exceeds the maximum order of the predictors and/or error terms, ARMAr-LASSO consistently outperforms the benchmark methods (see Section 4.1.1).
We list some important facts that capture how misspecification affects coefficient estimation and feature selection.
@@
(common AR(1) restriction). Suppose both predictors and error terms in model (1) admit a common AR(1) representation; that is, $x_{i,t}=\phi x_{i,t-1}+u_{i,t}$ and $\varepsilon_{t}=\phi\varepsilon_{t-1}+\omega_{t}$. In this case $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t}+\varepsilon_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}(\phi x_{i,t-1}+u_{i,t})+\phi\varepsilon_{t-1}+\omega_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}u_{i,t}+\phi y_{t-1}+\omega_{t}.$ Thus, under the common AR(1) restriction (also known as common factor restriction, Mizon1995), the working model (1) is equivalent to the true model (1) because of the decomposition of the AR(1) processes $\{\mathbf{x}_{t}\}$ and $\{\varepsilon_{t}\}$.
@@
The working model (1) coincides with the true model (1) under a common AR($p$) restriction; that is, when $x_{i,t}=\sum_{l=1}^{p}\phi_{l}x_{i,t-l}+u_{i,t}$ and $\varepsilon_{t}=\sum_{l=1}^{p}\phi_{l}\varepsilon_{t-l}+\omega_{t}$. In fact, it is easy to show that $y_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}x_{i,t}+\varepsilon_{t}=\sum_{i=1}^{n}\alpha_{i}^{*}u_{i,t}+\sum_{l=1}^{p}\phi_{l}y_{t-l}+\omega_{t}$ for any autoregressive order pp. Moreover, in this case $v_{t}=\omega_{t}$ and $E(v_{t}|\mathbf{w}_{t})=0$ – so we have unbiasedness and consistency also for the coefficients of the lags of $y_{t}$.
@@
(common factor). Suppose $x_{i,t}=\lambda_{i}f_{t}+z_{t},\ f_{t}=\phi_{f}f_{t-1}+\delta_{t},\ z_{i,t}=\phi_{i}z_{i,t-1}+\eta_{i,t}$ and $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t},$ where $\delta_{t},\eta_{it},\omega_{t}\sim i.i.d\ N(0,1).$ In this case, any $x_{it}$ is a sum of two independent AR(1) processes and, therefore, $x_{it}\sim ARMA(2,1)$ (Granger1976). Again, by Assumption 1, we have $E(v_{t}|\mathbf{u}_{t})=0$ and $E(v_{t}|y_{t-1})\neq 0$.
@@
Let $\mathbf{q}_{t}=(\mathbf{w}_{t}^{\prime},v_{t})^{\prime}$. To derive theoretical results for ARMAr-LASSO, we rely on the fact that, due to Assumption 1, $\mathbf{q}_{t}$ depends almost entirely on the “near epoch” of its shock. In particular, it is characterized as near-epoch dependent (NED) (refer to davidson1994 ch. 17 and ADAMEK2023 for details). NED is a very popular tool for modelling dependence in econometrics. It allows for cases where a variable’s behaviour is primarily governed by the recent history of explanatory variables or shock processes, potentially assumed to be mixing. davidson1994 shows that even if a variable is not mixing, its reliance on the near epoch of its shocks makes it suitable for applying limit theorems, particularly the mixingale property (see Supplement A.5 for details). The NED framework accommodates a wide range of models, including those that are misspecified as our working model (1). For instance, in Examples 2 and 3, $(\mathbf{w}_{t}^{\prime},v_{t})$ have a moving average representation with geometrically decaying coefficients, and are thus NED on $(\mathbf{u}_{t}^{\prime},\omega_{t})$ and $(\delta_{t},\boldsymbol{\eta}_{t}^{\prime},\omega_{t})$, respectively.
3.1 Estimation of ARMA residuals
In this section, we summarize the asymptotic properties of the Bayesian Information Criterion (BIC) and the Maximum Likelihood (ML) estimator for identifying ARMA orders and estimating model parameters, respectively.
[…]
Note that, under model misspecification, parameter estimates converge to pseudo‑true values and residuals converge to pseudo‑innovations rather than the true ones (see, e.g., potscher1991). These properties justify using the estimated coefficients and residuals as asymptotically valid approximations to their true counterparts.
3.2 Least Squares estimator applied to the Working Model (1)
To clarify the statistical properties of the coefficients penalized by ARMAr-LASSO, we first analyze the behavior of the corresponding Least Squares estimator applied to the ARMAr working model (ARMAr-LS). Consider the univariate model $y_{t}=\alpha x_{t-1}+\varepsilon_{t}$, where $x_{t}=\phi x_{t-1}+u_{t}$ and $\varepsilon_{t}=\phi_{\varepsilon}\varepsilon_{t-1}+\omega_{t}$, with $u_{t}$ and $\omega_{t}$ being serially uncorrelated innovations. Serial correlation is therefore present in both the predictor and the error term. The ARMAr transformation yields the working model $y_{t}=\alpha u_{t-1}+\phi_{y}y_{t-1}+v_{t}$, where the regressor $u_{t-1}$ is serially uncorrelated by construction, and the composite error $v_{t}$ collects the remaining dynamic components. Although $v_{t}$ is generally correlated with $y_{t-1}$, it is conditionally mean–independent of $u_{t-1}$ under Assumption 1. Consequently, the ARMAr-LS estimator satisfies $\widehat{\alpha}=\alpha+\frac{\sum_{t=1}^{T-1}u_{t}v_{t}}{\sum_{t=1}^{T-1}u_{t}^{2}}$, which immediately implies three key properties.
First, the estimator is unbiased, since $E(v_{t}|\mathbf{u})=0$ and therefore $E(\widehat{\alpha}|\mathbf{u})=\alpha$. Second, it is consistent, because $plim(\widehat{\alpha})=\alpha$ by exogeneity of $u_{t}$. Third, under the common AR(1) restriction $\phi=\phi_{\varepsilon}$, ARMAr-LS attains GLS efficiency, with $v_{t}=\omega_{t}$ and conditional variance $Var(\widehat{\alpha}|\mathbf{u})=\frac{\sigma_{\omega}^{2}}{\sum_{t=1}^{T-1}u_{t}^{2}}$.
These results are crucial for understanding the behavior of ARMAr-LASSO. In contrast to standard LASSO, which penalizes OLS coefficients that may be inefficient, biased, or inconsistent in the presence of serially correlated predictors Keele2005, ARMAr-LASSO penalizes coefficients that are already unbiased and consistent.
3.3 ARMAr-LASSO: Asymptotic Results
@@
This section is devoted to the asymptotic behaviour and feature selection consistency of the LASSO applied to working model (1), within the classical setting with $n$ fixed and $T\rightarrow\infty$. We will extend some known results to our context to demonstrate that the working model (1) retains the usual inferential and selection consistency properties, despite being a misspecification of the true model (1). Our results build upon Theorem 2 of Fu2000 and Theorem 1 of zhaoyu2006. In the classic asymptotic setting, the facts summarized in Section 3.1 allow us to derive properties directly for $u_{t}$ rather than for the estimate $\widehat{u}_{t}$. Let $\boldsymbol{\mu}_{vy}=\mathopen{}\mathclose{{\left(E(v_{t}y_{t-1}),\dots,E(v_{t}y_{t-p_{y}})}}\right)^{\prime}$ be the mean vector and $\boldsymbol{\Gamma}_{vy}$ the $p_{y}\times p_{y}$ covariance matrix of $\mathopen{}\mathclose{{\left(v_{t}y_{t-1},\dots,v_{t}y_{t-p_{y}}}}\right)$. The following theorem provides the asymptotic behaviour of the LASSO solution.
@@
(strong irrepresentable condition (zhaoyu2006)) Assuming $\widehat{\mathbf{C}}_{w}(11)$ is invertible, $|\widehat{\mathbf{C}}_{w}(21)(\widehat{\mathbf{C}}_{w}(11))^{-1}Sign(\boldsymbol{\beta}^{*}(1))|<\boldsymbol{1}-\varphi,$ where $\varphi\in(0,1)$ and the inequality holds element-wise.
@@
Let Assumptions 1 and 1 hold. If $\lambda\sqrt{T}\rightarrow\lambda_{0}\geq 0$, then the solution $\widehat{\boldsymbol{\beta}}$ of (2) is such that $P\mathopen{}\mathclose{{\left(Sign(\widehat{\boldsymbol{\beta}})=Sign(\boldsymbol{\beta}^{*})}}\right)\rightarrow 1.$
The theoretical results provided in this section show that under Assumptions 1 and 1, and as a consequence of Fact 1, ARMAr-LASSO guarantees consistent estimation, asymptotic normality, as well as consistent feature selection for the vector $\boldsymbol{\alpha}^{*}$.
Remark 3:
The working model (2) underlying ARMAr-LASSO constitutes a misspecification of the true data-generating process, which induces endogeneity due to lags of $y_{t}$. Theorems 1 and 2 are designed to show that, despite this endogeneity, ARMAr-LASSO preserves the classical properties of estimation and inference for the parameters of interest, namely the $\alpha$ coefficients associated with the ARMA residuals. Establishing these results is important to confirm that ARMAr-LASSO is a reliable tool for improving the estimation and forecasting performance of LASSO in settings where both predictors and errors exhibit serial correlation. Building on this foundation, we next turn to asymptotic results in a high-dimensional setting where both $n$ and $T$ diverge. This requires different conditions on the regularization parameter $\lambda$. Indeed, the optimal scaling of $\lambda$ depends on the setting: in high-dimensional asymptotics it typically follows $\sqrt{\log n/T}$, whereas in classical fixed-$n$ asymptotics a different scaling applies. Hence, distinct asymptotic regimes naturally imply different choices of $\lambda$ (see, e.g., Buhlmann2011; zhaoyu2006).
3.4 ARMAr-LASSO: Oracle Inequalities
In this section, we derive the oracle inequalities that provide bounds for the estimation and forecast errors of the ARMAr-LASSO. Here, we allow $n$ to grow as $T$ grows; that is, we pursue results in a framework where $n=n_{T}=O(T^{a})$ for some $a>0$. This condition serves as a broad upper bound and accommodates a variety of growth rates. For example, $n_{T}$ could grow as slowly as $log(T)$ or as fast as $T^{a}$. In this context, we replace the predictor vector $\mathbf{w}_{t}$ with $\widehat{\mathbf{w}}_{t}=(\widehat{u}_{1,t},\dots,\widehat{u}_{n_{T},t},y_{t-1},\dots,y_{t-p_{y}})^{\prime}$, where the $\widehat{u}$’s are obtained by employing BIC and ML as described in Section 3.1. We need the following Assumption, which bounds the unconditional moments of the predictors in the true model (1), and of $\widehat{\mathbf{w}}_{t}$ and $v_{t}$.